# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v2

@@ 3.3 Accounting for Nodes’ Knowledge of Injected Noise
We have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes. In practice, this likely has a small effect on the the DP guarantees. However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees. Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by allouahprivacy is scheme called SecLDP.
@@
For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm. 8, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. Similar correction can be carried out also in all the other cases. We do this correction in all our experiments.
3.4 Adaptive Compositions and Connection to Matrix-Mechanism Accounting
@@
The recent work by bellet2025unified shows that the linear systems perspective given in Theorem 8 can be equivalently analyzed via a so-called matrix mechanisms (kairouz2021practical; pillutla2025correlated). This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis. We next shortly discuss this connection. At a high level, both frameworks reduce an observer’s view to a linear image of a data-dependent signal plus correlated Gaussian noise.
Different Threat Scenarios.
@@
The matrix mechanism formulation by bellet2025unified allows analyzing private decentralized optimization methods that have a linear structure. Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
@@
  • The LDP approach where the adversary sees all messages (bellet2018personalized). This corresponds to the setting of Theorem 8 with $S=I_{n}$.
@@
  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (networkdp2). This corresponds to Theorem 8 with a selector matrix of a finite collection of nodes.
@@
  • SecLDP (allouahprivacy), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees. This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Thmeorem 8, see Section 3.3.
Privacy Analysis of Matrix Mechanisms.
@@
The matrix-mechanism analysis by bellet2025unified provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings. The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
@@
Definition 3.1 (Sensitivity under participation).
@@
Adaptive Analysis.
Notably, bellet2025unified extend the existing adaptive analysis of matrix mechanisms (denisov2022improved) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns. The crucial implication is that once the representation
@@
Matrix Mechanism Formulation of Gossip Averaging.
We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step. As shown in (Appendix A.4 bellet2025unified), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm. 8.
Equivalence of Sensitivity Bounds.
We also see from Thm. 10 that the sensitivity upper bound equals the upper bound given in our Thm. 8 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
@@
and the summation in the upper bound of Thm. 10 equals the upper bound given in our Thm. 8 in case the differing node participates in every round.
4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
We first need the following auxiliary lemma.
Lemma 11.
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic. Then there exist constants $M<\infty$ and $\rho\in(0,1)$ such that for all $t\geq 1$
$W^{t}\;=\;\Pi\;+\;R^{t},\qquad\Pi:=\mathbf{1}\,\pi^{\top},\qquad\|R^{t}\|\;\leq\;M\,\rho^{t},$
where $\pi^{\top}W=\pi^{\top}$, $\pi>0$, $\sum_{i}\pi_{i}=1$.
Remark 12.
A row-stochastic primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value < 1, and as the proof of Lemma 11 shows, $1-\rho$ equals the spectral gap $\lambda_{1}-\max_{i\geq 2}\left|\lambda_{i}\right|$.
Theorem 13.
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic with stationary distribution $\pi>0$, $\sum_{k=1}^{n}\pi_{k}=1$, $\pi^{\top}W=\pi^{\top}$. Let $H_{T}$ be the system matrix given in Lemma 7 with $S=e_{i}^{\top}$ (view of a single node, i.e., that of secure summation protocol) and define
as in Theorem 8. Then
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
Remark 14.
Notice that if $W$ is doubly stochastic, $\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}=\frac{1}{n}$ and we asymptotically have the same sensitivity as in case of a central aggregator.
5 Experiments
Illustration of the DP Bounds for Secure Summation Protocols. We first consider numerical evaluation of the DP bounds we obtain from Thm. 8.
First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$. We take the gossip matrix $W$ to be a row-stochastic matrix obtained from row-wise averaging. Second, we consider a network generated using a preferential-attachment model (barabasi1999emergence) with $n=200$ nodes, initialized from a fully connected core of 5 nodes. Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. The gossip matrix $W$ is obtained from row-wise averaging.
As Figure 1 illustrates, in case of the secure summation, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ for a randomly chosen $j$ grows asymptotically linearly w.r.t. $T$ as predicted by Thm. 13.
Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: gossip matrix of a preferential-attachment model with $n=200$ nodes, and the scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
6 Conclusions
We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that secure-summation-based averaging has asymptotic sensitivity comparable to central aggregation. Future work includes accelerating the linear-algebra routines needed to compute total sensitivities on large or sparse graphs and analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees.
Appendix A Proof of Lemma 6

### v3

@@ 3.3 Accounting for Nodes’ Knowledge of Injected Noise
We have so far assumed that the noise injected by the nodes belonging to the observing set $\mathcal{A}\in[n]$ contributes to the DP guarantees of the other nodes. We can obtain rigorous guarantees also in case they do not contribute, by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees. Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by Allouah et al. (2024) is scheme called SecLDP.
@@
For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm. 12, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. For a general adversarial set $\mathcal{A}$, we obtain a complete analogy of Lemma 10 and Thm. 12 with the matrix blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$.
3.4 Compute Efficient Upper Bound
We obtain a compute efficient upper bound for the sensitivity from a quadratic-form representation. Since $H_{T}$ has full row rank, $K_{T}:=H_{T}H_{T}^{\top}$ is invertible and for all $v\in\mathbb{R}^{m\cdot T\times 1}$
$\left\|H_{T}^{+}H_{T}v\right\|_{2}^{2}=(H_{T}v)^{\top}K_{T}^{-1}(H_{T}v).$
In particular, for $c\in\{-1,1\}^{T}$ and $v(c)=c\otimes e_{j}$, letting $b(c):=H_{T}v(c)$ we have
$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$
Moreover, we can construct a matrix $G_{T}\in\mathbb{R}^{mT\times T}$ such that $b(c)=G_{T}c.$ Hence, we can construct an $Tm\times Tm$ matrix $M_{T}:=G_{T}^{\top}K_{T}^{-1}G_{T}\succeq 0$, with which we get the efficiently computable upper bound
$\Delta_{T}(j\to\mathcal{A})^{2}$ $=\max_{c\in\{\pm 1\}^{T}}c^{\top}M_{T}c$
$\leq\max_{\|x\|_{2}^{2}=T}x^{\top}M_{T}x$
$=T\,\lambda_{\max}(M_{T}).$
We illustrate this bound numerically in Appendix F.
4 Asymptotic Sensitivity Growth Rate
Theorem 13.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic and primitive. Let $H_{T}$ be the system matrix given in Lemma 10 with $\mathcal{A}=\{i\}$, i.e., $S=e_{i}^{\top}$ for some $i\in[T]$ (single-node view) and define
as in Theorem 12. Then
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{1}{n}.$
Remark 14.
Notice that $(\Delta_{T}(j\to i))^{2}\sim T/n$ matches the sensitivity of the centralized case when each of the $n$ devices adds independent Gaussian noise with variance $\sigma^{2}$: their average has the output noise variance $n\sigma^{2}$ and as a result the DP analysis is equivalent to that of a Gaussian mechanism with noise scale $\sigma$ and squared sensitivity $T/n$ over a $T$-wise composition.
Using the proof technique of Thm. 13, we can also obtain the following sensitivity bound for the case the observing set $\mathcal{A}$ is of size $m$, $m>1$. Notice that Theorem 15 does not give a lower bound: as the proof indicates, the tight sensitivity becomes dependent on the gossip matrix $W$ and the observing set $\mathcal{A}$. Although the proof of Theorems 15 and 16 are analogous to the proof of Theorem 13 we state all these results separately for readability.
Theorem 15.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly-stochastic, and primitive. Fix a set of observing nodes $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$, and let $S_{\mathcal{A}}\in\mathbb{R}^{m\times n}$ be the row-selector matrix that extracts the coordinates indexed by $\mathcal{A}$ (i.e., $S_{\mathcal{A}}x=x_{\mathcal{A}}$). Let $H_{T}$ be the system matrix of Lemma 10 (Eq. (3.2)) constructed with observation matrix $S=S_{\mathcal{A}}$.
For any $j\in\{1,\dots,n\}$ define the sensitivity
$\Delta_{T}(j\to\mathcal{A}):=\max_{c\in\{\pm 1\}^{T}}\Bigl\|H_{T}^{+}H_{T}\,(c\otimes e_{j})\Bigr\|_{2}.$
Then, denoting $\rho:=\max\{\left|\lambda_{2}(W)\right|,\left|\lambda_{n}(W)\right|\}$, we have that for all $T\geq 1$,
$\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T+\frac{2}{(1-\rho)^{2}}\sqrt{\frac{m}{n}\,T}+\frac{2}{(1-\rho)^{4}}$
which further implies $\lim_{T\to\infty}\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T$.
Using the proof technique of Thm. 13, we also get an asymptotic upper bound for the case there are $m$ observing nodes and the noise injected by them does not contribute to the DP guarantees. We remark that the proof of Thm. 16 gives a similar explicit upper bound as in Thm. 15.
Theorem 16.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Fix an observing set $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$. Suppose $j\notin\mathcal{A}$ and suppose the noise terms injected by the nodes in $\mathcal{A}$ do not contribute to the DP guarantees. Denote by $\bar{\Delta}_{T}(j\to\mathcal{A})$ the sensitivity of Theorem 12 with the blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$. Then,
$\limsup_{T\to\infty}\frac{\bar{\Delta}_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n-m}.$
4.1 Experimental Illustration of the DP Bounds for Single Node View
We next consider a numerical illustration of the bound given by Theorem 13 in case a single node is observed. First, consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, i.e., each of the $n$ users is connected to each other with probability $p$. Second, consider a randomly drawn graph generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes. Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. This model produces a structure characterized by heavy-tailed degree distributions. In both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
As Figure 1 illustrates, in case only a single node $i$ is viewed, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ grows asymptotically linearly w.r.t. $T$ with the rate $1/n$ as predicted by Thm. 13. Here both $i$ and $j$ are chosen randomly.
Figure 1: The scaled sensitivity $(\Delta_{T}(j\to i))^{2}/T$ as a function of $T$ for a) randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
5 Utility Bound for Strongly Convex Losses
We next show how our sensitivity analysis can be plugged into standard convergence analysis of decentralized optimization (Koloskova et al., 2020) to obtain explicit privacy-utility bounds for private decentralized learning based on gossip averaging. For simplicity, we focus on the case of single-node view, however the results can be directly generalized to multi-node view using Theorems 15 and 16.
Setup. We consider decentralized minimization of the loss over $n$ nodes,
$f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$
over a connected undirected graph with a fixed symmetric, doubly-stochastic gossip matrix $W$. Assume each $f_{i}$ is $L$-smooth and $\mu$-strongly convex. Each node performs full-batch DP gradient descent, followed by gossip averaging:
$\theta_{i}^{t+\frac{1}{2}}$ $=\theta_{i}^{t}-\eta\,(\nabla f_{i}(\theta_{i}^{t})+\xi_{i}^{t}),$ (5.1)
$\theta_{i}^{t+1}$ $=\sum_{j=1}^{n}w_{ij}\,\theta_{j}^{t+\frac{1}{2}},$
where $\xi_{i}^{t}\sim\mathcal{N}(0,\sigma_{\mathrm{DP}}^{2}I_{d})$ are independent across $i,t$.
Following (Koloskova et al., 2020), we measure utility in terms of the average iterate $\bar{\theta}^{t}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}^{t}.$
Topology parameter (absolute spectral gap). Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be eigenvalues of $W$ and define the absolute spectral gap
$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$ (5.2)
In the decentralized SGD analysis by Koloskova et al. (2020), the effect of the network enters through a consensus-rate parameter $p$; for fixed symmetric $W$ one may take (see Lemma J.1 in Appendix)
$p=1-\gamma^{2}\;\;\;\text{and hence}\;\;\;\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right).$ (5.3)
Using Sensitivity Bound to Set DP Noise. Under the threat model of a single-node view $S=e_{i}^{\top}$, our Theorem 13 shows that the total sensitivity from node $j$ to observer $i$, $\Delta_{T}(j\to i))^{2}$ is asymptotically $T/n$ for doubly-stochastic $W$. We get an explicit upper bound from Thm. 15:
$(\Delta_{T}(j\to i))^{2}\;\leq\;\frac{T}{n}+\frac{2}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{2}{\gamma^{4}}.$ (5.4)
To guarantee $(\varepsilon,\delta)$-DP, we set the Gaussian noise as a function of $\varepsilon$ and $\delta$ via the standard condition (Dwork and Roth, 2014)
$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\,\sqrt{2\log(1.25/\delta)}}{\varepsilon}.$ (5.5)
Combining (5.5) with (5.4) gives the per-step DP noise scale
$\sigma_{\mathrm{DP}}^{2}\;\lesssim\;\frac{\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right).$ (5.6)
Gradient Noise for Convergence Bounds of (Koloskova et al., 2020). Because we use full-batch gradients, the only stochasticity is the injected DP noise, and the gradient-noise variances needed for the results of (Koloskova et al., 2020) are given by
$\sigma_{i}^{2}:=\mathbb{E}\|\xi_{i}^{t}\|_{2}^{2}=d\,\sigma_{\mathrm{DP}}^{2},$ (5.7)
which implies the mean noise parameter $\bar{\sigma}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}=d\,\sigma_{\mathrm{DP}}^{2}.$ Following Koloskova et al. (2020), we also define the heterogeneity measure
$\bar{\zeta}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\theta^{\star})\|_{2}^{2}.$
5.1 Utility Bound for DP Decentralized Learning
We can now directly apply the analysis by Koloskova et al. (2020) by using (5.6) and (5.7). We state the result in a form that the topology appears via $\gamma$ and privacy via $(\varepsilon,\delta)$. More details are given Appendix J.
Theorem 17 (Utility of Full-Batch DP-GD Training with Gossip Averaging).
Assume $f_{i}$ are $L$-smooth and $\mu$-strongly convex, and $W$ is symmetric, doubly-stochastic with absolute spectral gap $\gamma$. Run (5.1) for $T$ rounds with constant stepsize $\eta\simeq 1/L$ and Gaussian noise calibrated so that any node $j$ is $(\varepsilon,\delta)$-DP from the point of view of any node $i$, using (5.5) with $\Delta_{T}(j\to i)$ and the finite bound (5.4). Then the time-averaged excess risk of the average iterate satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{\theta}^{t})-f^{\star}\right]\;\leq$ $\widetilde{O}\!\bigg(\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n^{2}\,\mu}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{2}\sqrt{T}}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{4}T}$
$\quad\;+\;\frac{L\,\bar{\zeta}^{2}}{\mu^{2}\,\gamma^{2}\,T^{2}}\;+\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{\gamma}\exp\!\Big(-c\,\gamma\frac{\mu}{L}T\Big)\bigg),$
[…]
Proof on a High Level. Theorem 17 is obtained by: (i) viewing full-batch DP gradient descent with gossip averaging as decentralized SGD with gradient noise variance $\bar{\sigma}^{2}=d\sigma_{\mathrm{DP}}^{2}$, (ii) applying the strongly-convex convergence bound by Koloskova et al. (2020) with network parameter $p$ and substituting $p=\Theta(\gamma)$, and (iii) determining $\sigma_{\mathrm{DP}}$ via our single-node-view sensitivity given in Section 4.
Interpretation and Comparison to Centralized DP. The leading term in (17) is a privacy-induced error term $\Theta\!\left(\frac{d\log(1/\delta)}{\varepsilon^{2}n^{2}\mu}\right),$ which matches the scaling one has in the centralized setting (Bassily et al., 2014, 2019, 2020) (see also Sec. 4.2, Ponomareva et al., 2023). The remaining terms are transient and topology-dependent via the absolute spectral gap $\gamma$, vanishing as $T$ grows. Thus, the topology affects how quickly one reaches the privacy-utility baseline of the centralized setting.
6 Conclusions
We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that private gossip averaging gives asymptotically privacy guarantees comparable to central aggregation in case of a single-node view which further leads to utility results comparable to those of centralized private convex learning. We also analyzed the case of multiple-node view and SecLDP, the scenario where the noise injected by the viewing nodes does not contribute to the privacy guarantees. Future work includes analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees and privacy-utility trade-offs.
Appendix A Adaptive Compositions and Connection to Matrix-Mechanism Accounting
@@
The recent work by Bellet et al. (2025) shows that the linear systems perspective given in Theorem 12 can be equivalently analyzed via a so-called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025). This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis. We next shortly discuss this connection.
@@
Different Threat Scenarios. The matrix mechanism formulation by Bellet et al. (2025) allows analyzing private decentralized optimization methods that have a linear structure. Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
@@
  • The LDP approach where the adversary sees all messages (Bellet et al., 2018b). This corresponds to the setting of Theorem 12 with $S=I_{n}$.
@@
  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (Cyffers et al., 2022). This corresponds to Theorem 12 with a selector matrix of a finite collection of nodes.
@@
  • SecLDP (Allouah et al., 2024), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees. This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Theorem 12, see Section 3.3.
@@
Privacy Analysis of Matrix Mechanisms. The matrix-mechanism analysis by Bellet et al. (2025) provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings. The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
@@
Definition A.2 (Sensitivity under participation).
@@
Adaptive Analysis. Notably, Bellet et al. (2025) extend the existing adaptive analysis of matrix mechanisms (Denisov et al., 2022) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns. The crucial implication is that once the representation
@@
Matrix Mechanism Formulation of Gossip Averaging. We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step. As shown in (Appendix A.4 Bellet et al., 2025), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm. 12.
Equivalence of Sensitivity Bounds of Thm. 12 and of (Bellet et al., 2025).We also see from Thm. A.3 that the sensitivity upper bound equals the upper bound given in our Thm. 12 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
@@
and the summation in the upper bound of Thm. A.3 equals the upper bound given in our Thm. 12 in case the differing node participates in every round.
Appendix B Proof of Lemma 6